Determine the magnitude of F1 Express your answer to three significant figures and include the appropriate units. Suppose thaPLEASE SHOW ALL STEPS!!
Answer #1

ef F2 A F2 n:15 mn A cable has only one lension Sine foemulae Using T F2 sin (90+ B) (&tab) (+2) Sin an2 53. 13 Yan T Ces Sin36.3T =han B = 36. 87 O, = 40 Ao 36.87 53.13 T T Cos(367) Cos (36-87) 53. 1353.13) rom his T= Sin (s3.13+5 3.13) 6 25 N T = 6